. Consider the following information on the stock market in a small economy. (LO2)

Company Shares Outstanding Price, beginning of year Price, end of year

1 100 $100 $94

2 1,000 $20 $25

3 10,000 $3 $6

a. Compute a price-weighted stock price index for the beginning of the year and the end of the year. What is the percentage change?

b. Compute a value-weighted stock price index for the beginning of the year and the end of the year. What is the percentage change?
Answer #1

A) price weighted stock price index for the beginning of the year and the end of the year

Beginning of the year price = $100 + $20 +$3 = $123

end of the year price = $94 + $25 +$6 = $ 125

125-123 = 2/123*100

So the percentage change is 1.62%

B) value weighted stock price index for the beginning and end of the year

Beginning of the year value weighted stock price index =

(100* $100)+(1000* $20)+(10000* $30). = 10000+ 20000 + 30000. = $ 60000

At the end of the year the value is

(100* $94) + (1000*$25) + (10000* $6) = 9400+25000+60000 = $94400

94400-60000 =34400 /60000*100

So,the percentage change is 57.3%